OpenStudy (anonymous):

Please simplify (X^2+4x-21)/(x+7)

OpenStudy (anonymous):

\[(x^2+4x-21)/(x+7)\]

OpenStudy (anonymous):

\[{(x-3)(x+7)}/(x+7)=(x-3) so x=3\]

OpenStudy (anonymous):

@sstrin

OpenStudy (bradely):

X^2+4x-21 =x^2+7x-3x-21 =x(x+7)-3(x+7) =(x+7)(x-3) (X^2+4x-21)/(x+7)=(x+7)(x-3)/(x+7) =x-3 Source: http://www.mathskey.com/question2answer/